OpenStudy (anonymous):

Given a geometric sequence in the table below, create the explicit formula and list any restrictions to the domain. n an 1 −4 2 20 3 −100 an = −5(−4)n − 1 where n ≥ 1 an = −4(−5)n − 1 where n ≥ 1 an = −4(5)n − 1 where n ≥ −4 an = 5(−4)n − 1 where n ≥ −4 @D3xt3R

OpenStudy (anonymous):

We've got a sequence\[a_1=-4\\a_2=20\\a_3=-100\]What do we have in common?!\[a_1=-4\\a_2=(-4)*(-5)=20\\a_3=(-4)*(-5)*(-5)=(-4)*(-5)^2=-100\]

OpenStudy (anonymous):

? @D3xt3R

OpenStudy (anonymous):

You have to finish the exercise, how can you make a general term?!

OpenStudy (anonymous):

idk, i didnt learn this. im taking the class over online

OpenStudy (anonymous):

Think with me.\[a_1=-4\]\[a_2=(-4)*(-5)\]\[a_3=(-4)*(-5)^2\]Do you understand?!

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[a_1=-4\\a_2=20\\a_3=-100\]Do you understand it?!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What do you think, if I say the general term is\[a_n=a_1*(-5)^{n-1}\]??? Do you understant what I do?!

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

the terms in a geometric series are produced by multiplying b y a constant this is called the common ratio a simple example would be 1 , 2 , 4 , 8 where each term is multiplied by 2 to get the next one 1*2 = 2 2 * 2 = 4 4 * 2 = 8 etc here the common ratio is 2 common ratio is usually denoted by r and first term as a so we can write the sequence as a , ar, ar^2 , ar^3 note that the power on r is going up 1 at a time and the 3rd term has power 2, the fourth term has power 3 , and so on. so we can write the general nth term as a r^(n-1)

OpenStudy (cwrw238):

another convention is to write the first , second. third terms as a1, a2 , a3 etc

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

notice also that you can find the common ratio r by dividing the first term by the first, or third term by the second , 4th by the third etc because ar/a = r and ar^2 / ar = r etc so to find r in the sequence in the question we can divide 2nd by the first - that is 20 / -4 = -5 -100 / 50 = -5 so common ratio r = -5

OpenStudy (cwrw238):

so to find an - that id the nth term of the series in the wuestion we have the general formula an = ar^(n-1) now a = -4 and r = -5 so an = -4(-5)^(n-1) now n being the sequance number of the term must be 1 or greater so we write n>+1

OpenStudy (cwrw238):

n >= 1 so the 2nd option is the correct one.

OpenStudy (cwrw238):

Formula to commit to memory for a GS is nth term = ar^(n-1) or if you like nth term = a1r^(n-1) a or a1 is the first term

OpenStudy (anonymous):

thank you